# Campana points on wonderful compactifications

### v3

@@ 5.2. Good primes
Next, we view $\beta$ in $\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{T}}^{\prime}_{F_{v}})$. The restriction of $\beta$ to ${\mathsf{S}}^{\prime}_{v}$ is an element of $\theta_{v}\in\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{S}}^{\prime}_{v})$. For $\theta\in\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{S}}^{\prime}_{v})$, let
$l_{v}(\eta)=\#\{\beta\in\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{T}}^{\prime}_{F_{v}})\ \textrm{with}\ \beta|_{{\mathsf{S}}^{\prime}_{v}}=\theta\}.$
@@ 5.3. Bad primes
${\operatorname{Re}s_{\alpha}>\kappa_{\alpha}\text{ for all }\alpha\in\mathcal{A}}.\qed$
@@ 5.4. Places in $S$
$|\prod_{\alpha\in\mathcal{A}}L_{v}(s_{\alpha}-\kappa_{\alpha},\chi_{\alpha})^{-1}\mathcal{J}_{v}({\mathbf{s}})|\leq C_{v}({\epsilon})\quad\text{ and }\quad|\prod_{\alpha\in\mathcal{A}}L_{v}(s_{\alpha}-\kappa_{\alpha},\chi_{\alpha})^{-1}\mathcal{J}_{v,\partial}({\mathbf{s}})|\leq C_{v}(\partial,{\epsilon}),$
@@ 6. Height integrals II: Infinite-dimensional representations
In this section we will prove that, as in the case of rational points, the remaining terms in the spectral decomposition do not contribute to the rightmost pole. Here we adapt the proof of [LTBT20, Theorem 4.10] . In order to do this we just need to prove an analogue of [STBT07, Corollary 7.4] in our context. Throughout this section, we assume that the quasi-split inner form of ${\mathsf{G}}$ has rank $\geq 2$ The results hold even if $(X,D_{\epsilon})$ is not klt, though our bounds are only useful in the klt case (see Theorem 6.6).
@@
is holomorphic on $\operatorname{Re}s_{\alpha}>\kappa_{\alpha}+1/m_{\alpha}-1/3n$ for all $\alpha\in\mathcal{A}$. Here $n$ is as in Proposition 6.1. For all ${\epsilon}>0$ and all compact subsets $K\subset\{{\bf s}\in\mathrm{Pic}(X)_{\mathbb{C}}\,|\,\operatorname{Re}s_{\alpha}>\kappa_{\alpha}+1/m_{\alpha}-1/3n+\epsilon\ \text{for all}\ \alpha\in\mathcal{A}\}$, there is a constant $C({\epsilon},K)$, independent of $\pi$, such that
@@
$S^{\flat}(\mathbb{s})=\sum_{\chi\in\mathfrak{X}}\sum_{P}n(A_{P})^{-1}\int_{\Pi(M_{P})}(\sum_{\phi\in B_{P}(\pi)_{\chi}}E(e,\phi)\int_{G(\mathbb{A}_{F})}\delta_{\epsilon}(g)\mathsf{H}(\mathbb{s},g)^{-1}\overline{E(g,\phi)}\,\mathrm{d}g)\,\mathrm{d}\pi_{p}.$
@@
The proof is identical to the proof of Theorem 4.10 of [LTBT20]. The key ingredient is the analogue of [LTBT20, Theorem 4.9] in our setting and this theorem, similar to [STBT07, Corollary 7.4], follows from [LTBT20, Lemma 4.7], Theorem 6.2, Proposition 6.3, and Corollary 6.5. Here for the convenience of the reader we provide a very brief outline of the proof. For each archimedean place $v$, let ${\mathsf{K}}_{v}$ be a maximal compact subgroup as in Proposition 2.1. Let ${\mathsf{G}}_{\infty}=\prod_{v\text{ arch.}}{\mathsf{G}}(F_{v})$ and ${\mathsf{K}}_{\infty}=\prod_{v\text{ arch.}}{\mathsf{K}}_{v}$. For each finite dimensional smooth representation $W$ of ${\mathsf{K}}_{\infty}$, let $\mathrm{tr}_{W}$ be its trace function extended to ${\mathsf{G}}_{\infty}$ by defining it to be zero outside ${\mathsf{K}}_{\infty}$. We let $\xi_{W}=(\mathrm{dim}\,W)\cdot\mathrm{tr}_{W}$. For each non-archimedean place $v$, pick ${\mathsf{K}}_{v}$ as in Proposition 4.5, and let $\xi_{v}=\mathrm{vol}\,({\mathsf{K}}_{v})^{-1}ch_{{\mathsf{K}}_{v}}$ where here $ch_{{\mathsf{K}}_{v}}$ is the characteristic function of the set ${\mathsf{K}}_{v}$. Let $\xi_{f}=\otimes_{v\text{ non-arch}}\xi_{v}$. By a theorem of Harish-Chandra [War72, Prop. 4.4.3.2.], $\sum_{W\in\hat{\mathsf{K}}_{\infty}}(\xi_{f}\otimes\xi_{W})*(\delta_{\epsilon}(g){\mathsf{H}}(\mathbf{s},g)^{-1})$ converges in the topology of $C^{\infty}({\mathsf{G}}({\mathbb{A}}))$ to $\delta_{\epsilon}(g){\mathsf{H}}(g)$. Now we have
@@
after a change of variables. The key point is that $(\xi_{f}\otimes\xi_{W})*E(.,\phi)$, modulo some harmless functions, is the product $\prod_{v\not\in S}\varphi_{\pi_{v}}$ where here $\pi=\otimes_{v}^{\prime}\pi_{v}$ is the automorphic representation for $E(.,\phi)$ and $\varphi_{\pi_{v}}$ is the normalized spherical function associated to the representation $\pi_{v}$ (for details see Lemma 4.7 of [LTBT20]). Now that we have spherical functions we can use the bounds obtained in Theorem 6.2 for $v\not\in S$ and Proposition 6.3, and Corollary 6.5 for $v\in S$. Once we use these bounds, the convergence of the resulting sum is a result of Müller [M9̈8, Corollary 0.3]; for details see the proof of Theorem 4.10 of [LTBT20].
In the klt case we have $01/m_{\alpha}>$ 0, thus $\max\{\kappa_{\alpha}+1/m_{\alpha}-1/3n,\kappa_{\alpha}\}<\kappa_{\alpha}+1/m_{\alpha}$. ∎
@@ 7. The leading pole
$\delta^{\prime}_{\epsilon,v}(P_{v})=1\quad\iff\quad\begin{array}[]{ll}&\text{ for all $\alpha\in\mathcal{A}(L)$ we have $m_{\alpha}\mid n_{v}(\mathcal{D}_{\alpha},P)$}\\ &\text{ and for all $\alpha\notin\mathcal{A}(L)$ we have $n_{v}(\mathcal{D}_{\alpha},P_{v})=0,$}\end{array}$
@@ 8. Conjecture for the leading constant
In [PSTVA21, Section 3.3] a conjecture was put forward for the leading constant for counting Campana points of bounded height. This conjecture was shown to be incorrect independently by Streeter [Str22] and Shute [Shu22]. The problematic part of the conjecture is the factor #$\#\mathrm{H}^{1}(\Gamma,\operatorname{Pic}\bar{U})$ proposed as a replacement for the factor $\beta(X)=\#\mathrm{H}^{1}(\Gamma,\operatorname{Pic}\bar{X})$ in Manin’s conjecture. We suggest a fix which is compatible with Theorem 7.1. It is somewhat subtle as it involves taking a sum over integrals with respect to the Brauer–Manin pairing and Tamagawa measure, since the resulting Brauer group can be infinite (modulo constants).
@@ 8.1. Set-up
We follow closely the set-up from [PSTVA21, Section 3.3]. Let $X$ be a smooth projective variety over a number field $F$ and let $(X,D_{\epsilon})$ be a smooth klt Campana orbifold with $D_{\epsilon}=\sum_{\alpha\in\mathcal{A}}\epsilon_{\alpha}D_{\alpha}$, for some $D_{\alpha}$. For each $\alpha\in\mathcal{A}$, we fix an adelic metrization for $\mathcal{O}(D_{\alpha})$ and denote its adelically metrized divisor as $\mathcal{D}_{\alpha}$. Let $\mathcal{L}$ be an adelically metrised log adjoint rigid big and nef divisor so that $a((X,D_{\epsilon}),L)L+K_{X}+D_{\epsilon}$ is $\mathbb{Q}$-linearly equivalent to an effective $\mathbb{Q}$-divisor $E$ on $X$ which has Iitaka dimension 00. We assume that the support of $E$ is contained in the support of $D$. We also fix the adelic metrization for the canonical divisor $K_{X}$ which is induced by adelic metrizations of $\mathcal{L}$ and $\mathcal{D}_{\alpha}$. Our assumptions imply that if we let $U=X\setminus\mathrm{Supp}(E)$ denote the complement of the support of $E$ in $X$, then we have $k[U]^{\times}=k^{\times}$. Indeed, if there is a non-constant and non-vanishing regular function on $U$, then it defines a non-trivial linear equivalence relation among irreducible components of $E$. However, this contradicts the fact that the Iitaka dimension of $E$ is 00.
@@ 8.2. Campana Brauer group
$\operatorname{Br}((X,D_{\epsilon}),L)=\{\beta\in\operatorname{Br}X^{\circ}:m_{D}\operatorname{\partial}_{D}(\beta)=0\,\,\forall D\subseteq U\setminus X^{\circ}\}.$
Here $D$ is a divisor occurring as an irreducible component of $U\setminus X^{\circ}$ and $\operatorname{\partial}_{D}:\operatorname{Br}X^{\circ}\to{\mathsf{H}}^{1}(\kappa(D),{\mathbb{Q}}/{\mathbb{Z}})$ denotes the residue map at $D$. This group (without reference to $L$) first appeared in the context of semi-integral points in [MNS22, Def. 3.14]. We define the algebraic part by $\operatorname{Br}_{1}((X,D_{\epsilon}),L)=\operatorname{Br}((X,D_{\epsilon}),L)\cap\operatorname{Br}_{1}X^{\circ}$. A particularly important case is the log anticanonical divisor $\operatorname{Br}((X,D_{\epsilon}),-K_{X}-D_{\epsilon})$, where one takes all boundary divisors $D_{\epsilon}$ in the definition since $U=X$; we call this simply the Campana Brauer group $\operatorname{Br}(X,D_{\epsilon})$ of the pair $(X,D_{\epsilon})$. The group $\operatorname{Br}(X,D_{\epsilon})/\operatorname{Br}{\mathbb{Q}}$ can be infinite even if $(X,D_{\epsilon})$ is log Fano. This complicates matters greatly and reflects the fact that in many examples, the orbifold fundamental group of $(X,D_{\epsilon})$, in the sense of [Cam11, Def. 11.1], can be non-trivial.
@@
Consider the pair $(\mathbb{P}^{1},(1/2)0+(1/2)\infty)$ over $\mathbb{Q}$. Then the Campana Brauer group is generated by $\operatorname{Br}\mathbb{Q}$ and all the quaternion algebras $(t,a)$ where $t$ is a coordinate function and $a\in{\mathbb{Q}}^{\times}$. These are all distinct as $a$ ranges over representatives of ×/×${\mathbb{Q}}^{\times}/{\mathbb{Q}}^{\times 2}$.
@@ 8.4. Verification for the wonderful compactification
$0\to M\to\operatorname{Pic}(U)\to\operatorname{Pic}({\mathsf{G}})\to 0$
@@
×$\times\lim_{s\to a}\int_{{\mathsf{G}}(\mathbb{A}_{F})}\prod_{\begin{subarray}{c}\alpha\in\mathcal{A}(L)\\ v\in\Omega_{F}\end{subarray}}\frac{1}{\zeta_{F_{\alpha},v}(m_{\alpha}(s\lambda_{\alpha}-\kappa_{\alpha}))}\delta_{\epsilon}(g)\mathsf{H}(sL,g)^{-1}\chi(g)\mathrm{d}g$
@@ 8.6.2. Local densities
therefore the integral is 00. Now assume that $p\nmid d$. Then $(d,t)_{p}=\left(\frac{d}{p}\right)^{v_{p}(t)}$. Using that the region of integration has measure $p^{-n}-p^{-n-1}$ completes the proof. ∎
@@
#$\#\{(a,c)\in{\mathbb{F}}_{p}^{2}:a^{2}+a=c^{2}\}=\sum_{a\in{\mathbb{F}}_{p}}\left(1+\left(\frac{a(1+a)}{p}\right)\right).$
However this is the number of ${\mathbb{F}}_{p}$-points on a smooth affine conic with two rational points at infinity, which is thus $p-1$. The result follows. ∎
@@
=$=\begin{cases}\left(\left(\frac{-d_{0}d_{1}}{p}\right)+\left(\frac{-d_{0}d_{2}}{p}\right)+\left(\frac{-d_{1}d_{2}}{p}\right)\right)p^{-3/2},\quad&p\mid d_{0}d_{1}d_{2},\\ 1+p^{-1}+\left(\left(\frac{d_{0}d_{1}}{p}\right)+\left(\frac{d_{0}d_{2}}{p}\right)+\left(\frac{d_{1}d_{2}}{p}\right)\right)p^{-3/2},\quad&p\nmid d_{0}d_{1}d_{2}.\end{cases}$
@@
The first step of our proof is the same as in the proof of [Shu22, Lem. 3.1]. We use the parametrisation $t\mapsto(t:1:-1-t)$. With respect to this, the $3$ marked points are $t=0,-1$ and $\infty$. As in [Shu22, (3.10)] we have that the integral equals
@@
where the symbol is the Legendre symbol. If $p\nmid d_{0}d_{1}d_{2}$ this equals $1-2/p$.
@@
Near to 00: Firstly assume that $v_{p}(t)\geq 2$ so that we are close to 00. Here $|1+t|_{p}=\max\{|t|_{p},1|\}=1$ and $(d_{i},-(1+t))_{p}=(d_{i},-1)_{p}$. Hence (8.9) becomes
@@
By Lemma 8.9 this is 00 for $p\mid d_{1}d_{2}$. If $p\nmid d_{1}d_{2}$ then $(d_{1},-1)_{p}=1$, hence we obtain that (8.10) equals $(d_{0},-1)_{p}$ times
@@
which, as in (8.10), is 00 for $p\mid d_{0}d_{2}$ and $(d_{1},-1)_{p}(p^{-1}+\left(\frac{d_{0}d_{2}}{p}\right)p^{-3/2})$ otherwise.
Near to $-1$: Finally close to $-1$, we have the integral
@@
We make the change of variables $u=-(1+t)$ to obtain
@@
As in (8.10), this equals 00 for $p\mid d_{0}d_{1}$ and $(d_{2},-1)_{p}(p^{-1}+\left(\frac{d_{0}d_{1}}{p}\right)p^{-3/2})$ otherwise. Combining all cases together gives the statement of the lemma. ∎
@@
×$\times\sum_{\begin{subarray}{c}\mathbf{y}\end{subarray}}\frac{\mu^{2}(d_{0}d_{1}d_{2}y_{0}y_{1}y_{2})\gamma(y_{0}y_{1}y_{2})}{(y_{0}y_{1}y_{2})^{3/2}}\left(\frac{y_{1}y_{2}}{d_{0}}\right)\left(\frac{y_{0}y_{2}}{d_{1}}\right)\left(\frac{y_{0}y_{1}}{d_{2}}\right).$

### v4

@@ 5.2. Good primes
Next, we view $\beta$ in $\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{T}}^{\prime}_{F_{v}})$. The restriction of $\beta$ to ${\mathsf{S}}^{\prime}_{v}$ is an element $\theta_{v}\in\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{S}}^{\prime}_{v})$. For $\theta\in\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{S}}^{\prime}_{v})$, let
$l_{v}(\theta)=\#\{\beta\in\Delta({\mathsf{G}}^{\prime}_{F_{v}},{\mathsf{T}}^{\prime}_{F_{v}})\ \textrm{with}\ \beta|_{{\mathsf{S}}^{\prime}_{v}}=\theta\}.$
@@ 5.3. Bad primes
$\operatorname{Re}s_{\alpha}>\kappa_{\alpha}\text{ for all }\alpha\in\mathcal{A}$$\qed$
@@ 5.4. Places in $S$
[…]
@@ 6. Height integrals II: Infinite-dimensional representations
In this section we will prove that, as in the case of rational points, the remaining terms in the spectral decomposition do not contribute to the rightmost pole. Here we adapt the proof of [LTBT20, Theorem 4.10]. In order to do this we just need to prove an analogue of [STBT07, Corollary 7.4] in our context. Throughout this section, we assume that the quasi-split inner form of $\mathsf{G}$} has rank $\geq 2$. The results hold even if $(X,D_{\epsilon})$ is not klt, though our bounds are only useful in the klt case (see Theorem 6.6).
@@
is holomorphic on $\operatorname{Re}s_{\alpha}>\kappa_{\alpha}+1/m_{\alpha}-1/3n$ for all $\alpha\in\mathcal{A}$. Here $n$ is as in Proposition 6.1. For all ${\epsilon}>0$ and all compact subsets
$K\subset\{{\bf s}\in\mathrm{Pic}(X)_{\mathbb{C}}\,|\,\operatorname{Re}s_{\alpha}>\kappa_{\alpha}+1/m_{\alpha}-1/3n+\epsilon\ \text{for all}\ \alpha\in\mathcal{A}\},$
there is a constant $C({\epsilon},K)$, independent of $\pi$, such that
@@
$S^{\flat}(\mathbb{s})=\sum_{\chi\in\mathfrak{X}}\sum_{P}n(A_{P})^{-1}\int_{\Pi(M_{P})}\left(\sum_{\phi\in B_{P}(\pi)_{\chi}}E(e,\phi)\int_{G(\mathbb{A}_{F})}\delta_{\epsilon}(g)\mathsf{H}(\mathbb{s},g)^{-1}\overline{E(g,\phi)}\,\mathrm{d}g\right)\,\mathrm{d}\pi_{p}.$
@@
[…]
@@
after a change of variables. The key point is that $(\xi_{f}\otimes\xi_{W})*E(.,\phi)$, modulo some harmless functions, is the product $\prod_{v\not\in S}\varphi_{\pi_{v}}$ where here $\pi=\otimes_{v}^{\prime}\pi_{v}$ is the automorphic representation for $E(.,\phi)$ and $\varphi_{\pi_{v}}$ is the normalized spherical function associated to the representation $\pi_{v}$ (for details see Lemma 4.7 of [LTBT20]). Now that we have spherical functions we can use the bounds obtained in Theorem 6.2 for $v\not\in S$ and Proposition 6.3, and Corollary 6.5 for $v\in S$. Once we use these bounds, the convergence of the resulting sum is a result of Müller [Mü98, Corollary 0.3]; for details see the proof of Theorem 4.10 of [LTBT20].
In the klt case we have $1/m_{\alpha}>0$, thus $\max\{\kappa_{\alpha}+1/m_{\alpha}-1/3n,\kappa_{\alpha}\}<\kappa_{\alpha}+1/m_{\alpha}$. ∎
@@ 7. The leading pole
[…]
@@ 8. Conjecture for the leading constant
In [PSTVA21, Section 3.3] a conjecture was put forward for the leading constant for counting Campana points of bounded height. This conjecture was shown to be incorrect independently by Streeter [Str22] and Shute [Shu22]. The problematic part of the conjecture is the factor $\mathrm{H}^{1}(\Gamma_{F},\operatorname{Pic}\bar{U})$ proposed as a replacement for the factor $\beta(X)=\#\mathrm{H}^{1}(\Gamma_{F},\operatorname{Pic}\bar{X})$ in Manin’s conjecture. We suggest a fix which is compatible with Theorem 7.1. It is somewhat subtle as it involves taking a sum over integrals with respect to the Brauer–Manin pairing and Tamagawa measure, since the resulting Brauer group can be infinite (modulo constants).
@@ 8.1. Set-up
We follow closely the set-up from [PSTVA21, Section 3.3]. Let $X$ be a smooth projective variety over a number field $F$ and let $(X,D_{\epsilon})$ be a smooth klt Campana orbifold with $D_{\epsilon}=\sum_{\alpha\in\mathcal{A}}\epsilon_{\alpha}D_{\alpha}$, for some $D_{\alpha}$. For each $\alpha\in\mathcal{A}$, we fix an adelic metrization for $\mathcal{O}(D_{\alpha})$ and denote its adelically metrized divisor as $\mathcal{D}_{\alpha}$. Let $\mathcal{L}$ be an adelically metrised log adjoint rigid big and nef divisor so that $a((X,D_{\epsilon}),L)L+K_{X}+D_{\epsilon}$ is $\mathbb{Q}$-linearly equivalent to an effective $\mathbb{Q}$-divisor $E$ on $X$ which has Iitaka dimension 0. We assume that the support of $E$ is contained in the support of $D$. We also fix the adelic metrization for the canonical divisor $K_{X}$ which is induced by adelic metrizations of $\mathcal{L}$ and $\mathcal{D}_{\alpha}$. Our assumptions imply that if we let $U=X\setminus\mathrm{Supp}(E)$ denote the complement of the support of $E$ in $X$, then we have $k[U]^{\times}=k^{\times}$. Indeed, if there is a non-constant and non-vanishing regular function on $U$, then it defines a non-trivial linear equivalence relation among irreducible components of $E$. However, this contradicts the fact that the Iitaka dimension of $E$ is 0.
@@ 8.2. Campana Brauer group
$\operatorname{Br}((X,D_{\epsilon}),L)=\{\beta\in\operatorname{Br}X^{\circ}:m_{\alpha}\operatorname{\partial}_{D_{\alpha}}(\beta)=0\,\,\forall D_{\alpha}\subseteq U\setminus X^{\circ}\}.$
Here $D_{\alpha}$ is a divisor occurring as an irreducible component of $U\setminus X^{\circ}$ and $\operatorname{\partial}_{D}:\operatorname{Br}X^{\circ}\to{\mathsf{H}}^{1}(\kappa(D),{\mathbb{Q}}/{\mathbb{Z}})$ denotes the residue map at $D$. This group (without reference to $L$) first appeared in the context of semi-integral points in [MNS22, Def. 3.14]. We define the algebraic part by $\operatorname{Br}_{1}((X,D_{\epsilon}),L)=\operatorname{Br}((X,D_{\epsilon}),L)\cap\operatorname{Br}_{1}X^{\circ}$. A particularly important case is the log anticanonical divisor $\operatorname{Br}((X,D_{\epsilon}),-K_{X}-D_{\epsilon})$, where one takes all boundary divisors $D_{\epsilon}$ in the definition since $U=X$; we call this simply the Campana Brauer group $\operatorname{Br}(X,D_{\epsilon})$ of the pair $(X,D_{\epsilon})$. The group $\operatorname{Br}(X,D_{\epsilon})/\operatorname{Br}{\mathbb{Q}}$ can be infinite even if $(X,D_{\epsilon})$ is log Fano. This complicates matters greatly and reflects the fact that in many examples, the orbifold fundamental group of $(X,D_{\epsilon})$, in the sense of [Cam11, Def. 11.1], can be non-trivial.
@@
Consider the pair $(\mathbb{P}^{1},(1/2)0+(1/2)\infty)$ over $\mathbb{Q}$. Then the Campana Brauer group is generated by $\operatorname{Br}\mathbb{Q}$ and all the quaternion algebras $(t,a)$ where $t$ is a coordinate function and $a\in{\mathbb{Q}}^{\times}$. These are all distinct as $a$ ranges over representatives of ${\mathbb{Q}}^{\times}/{\mathbb{Q}}^{\times 2}$.
@@ 8.4. Verification for the wonderful compactification
$0\to M\to\operatorname{Pic}(U)\to\operatorname{Pic}({\mathsf{G}})\to 0.$
@@
$\times\lim_{s\to a}\int_{{\mathsf{G}}(\mathbb{A}_{F})}\prod_{\begin{subarray}{c}\alpha\in\mathcal{A}(L)\\ v\in\Omega_{F}\end{subarray}}\frac{1}{\zeta_{F_{\alpha},v}(m_{\alpha}(s\lambda_{\alpha}-\kappa_{\alpha}))}\delta_{\epsilon}(g)\mathsf{H}(sL,g)^{-1}\chi(g)\mathrm{d}g$
@@ 8.6.2. Local densities
therefore the integral is 0. Now assume that $p\nmid d$. Then $(d,t)_{p}=\left(\frac{d}{p}\right)^{v_{p}(t)}$. Using that the region of integration has measure $p^{-n}-p^{-n-1}$ completes the proof. ∎
@@
$\{(a,c)\in{\mathbb{F}}_{p}^{2}:a^{2}+a=c^{2}\}=\sum_{a\in{\mathbb{F}}_{p}}\left(1+\left(\frac{a(1+a)}{p}\right)\right).$
However this is the number of ${\mathbb{F}}_{p}$-points on a smooth affine conic with two rational points at infinity, which is thus p1p-1. The result follows. ∎
@@
$=\begin{cases}\left(\left(\frac{-d_{0}d_{1}}{p}\right)+\left(\frac{-d_{0}d_{2}}{p}\right)+\left(\frac{-d_{1}d_{2}}{p}\right)\right)p^{-3/2},\quad&p\mid d_{0}d_{1}d_{2},\\ 1+p^{-1}+\left(\left(\frac{d_{0}d_{1}}{p}\right)+\left(\frac{d_{0}d_{2}}{p}\right)+\left(\frac{d_{1}d_{2}}{p}\right)\right)p^{-3/2},\quad&p\nmid d_{0}d_{1}d_{2}.\end{cases}$
@@
The first step of our proof is the same as in the proof of [Shu22, Lem. 3.1]. We use the parametrisation $t\mapsto(t:1:-1-t)$. With respect to this, the $3$ marked points are t=0,1t=0,-1 and $\infty$. As in [Shu22, (3.10)] we have that the integral equals
@@
where the symbol is the Legendre symbol. If $p\nmid d_{0}d_{1}d_{2}$ this equals 12/p1-2/p.
@@
Near to 0: Firstly assume that $v_{p}(t)\geq 2$ so that we are close to 0. Here $|1+t|_{p}=\max\{|t|_{p},1|\}=1$ and $(d_{i},-(1+t))_{p}=(d_{i},-1)_{p}$. Hence (8.9) becomes
@@
By Lemma 8.9 this is 0 for $p\mid d_{1}d_{2}$. If $p\nmid d_{1}d_{2}$ then $(d_{1},-1)_{p}=1$, hence we obtain that (8.10) equals $(d_{0},-1)_{p}$ times
@@
which, as in (8.10), is 0 for $p\mid d_{0}d_{2}$ and $(d_{1},-1)_{p}(p^{-1}+\left(\frac{d_{0}d_{2}}{p}\right)p^{-3/2})$ otherwise.
Near to 1-1: Finally close to 1-1, we have the integral
@@
We make the change of variables u=(1+t)u=-(1+t) to obtain
@@
As in (8.10), this equals 0 for $p\mid d_{0}d_{1}$ and $(d_{2},-1)_{p}(p^{-1}+\left(\frac{d_{0}d_{1}}{p}\right)p^{-3/2})$ otherwise. Combining all cases together gives the statement of the lemma. ∎
@@
$\times\sum_{\begin{subarray}{c}\mathbf{y}\end{subarray}}\frac{\mu^{2}(d_{0}d_{1}d_{2}y_{0}y_{1}y_{2})\gamma(y_{0}y_{1}y_{2})}{(y_{0}y_{1}y_{2})^{3/2}}\left(\frac{y_{1}y_{2}}{d_{0}}\right)\left(\frac{y_{0}y_{2}}{d_{1}}\right)\left(\frac{y_{0}y_{1}}{d_{2}}\right).$